OpenStudy (anonymous):

A car weighing 15680 N and moving at 20.0 /s is acted upon by a 6.50 x 10^2N force until it is brought to a halt a. what is the car's mass b. What is its initial momentum? c. What is the change in the car's momentum? d. How long does the braking force act on the car to bring it to a halt?

random231 (random231):

a.15680/10(9.8 would be more exact) b.1568*20=31360 c.final momentum - initial momentum=0-31360 d. we know f=ma a=f/m so here retardation is(a)=(6.50*10^2)/1568 use v=u-at v=0 u=20 a=(6.50*10^2)/1568